OpenStudy (tyrinzoebrown):

I need help, Solve the equation 13+w/7=-18 showing all needed steps

OpenStudy (anonymous):

13 + x / 7 = -18 13 + x = -18 * 7 13 + x = -126 x = -126-13 = -139

OpenStudy (anonymous):

w is x , btw

OpenStudy (tyrinzoebrown):

oh my God thank you so much!

OpenStudy (anonymous):

welcome